A company has just paid its annual dividend of $2.64 yesterday, and it is unlikely to change the amount paid out in future years. If the required rate of return is 19 percent p.a., what is the share worth today? (to the nearest cent; don’t include $ sign)
Answer #1

Dividend Paid out each year = D = $2.64

Rate of return = r = 19%

Present Value of the stock = D/(1+r) + D/(1+r)2 + .....

= D/r

= 2.64/0.19

= $13.89